# Agentic Large Language Models, a survey

### v2

@@ 2.4.2 Artificial General Intelligence
The work in this section, and especially the work on self reflection, connects to research on artificial general intelligence, in the scientific tradition of artificial intelligence that created strong narrow intelligence in backgammon (Tesauro, 1994), chess (Hsu, 2022; Müller and Schaeffer, 2018), and go (Silver et al., 2016). This tradition views intelligence as a competitive, individualistic, reasoning problem (Plaat, 2020). The benefits and risks related to super-intelligence and singularities are actively debated (Bostrom, 1998; Kurzweil, 2022), raising ethical and philosophical questions (Dennett, 2017). Here, intelligence is regarded as a feature of individuals. Most visions of super intelligence assume that the artificial agent has the ability to use tools and to function in a social environment, something that humans do easily. However, human-unique parts of intelligence emerge in social contexts and depend on constant interaction with others, both evolutionarily and developmentally. We will see work that focuses on social interaction by artificial agents in later sections.
2.4.3 Interpretability
@@
How do LLMs really work on the inside? Opening up the black box of neural connectionist architectures is an important topic of research. We wish to understand how the billions of neurons embed representations, how they reason, and how they come to conclusions. Explainable AI provides different methods to do so (Minh et al., 2022; Rios et al., 2020; Van Stein et al., 2022; Selvaraju et al., 2017; Ali et al., 2022).
@@
Static methods from the symbolic tradition have been quite successful (Molnar, 2020). Methods exist to relate how input pictures map to output classes, for example using feature maps (Ren et al., 2016; Kohonen, 1982; Redmon et al., 2016). Counterfactual analysis (Karimi et al., 2020; Huang et al., 2024b), LIME (Ribeiro et al., 2016), and SHAP (Lundberg and Lee, 2017) help understand how inputs map to outputs for structured data. Distillation methods can map neural networks to decision trees (Hinton et al., 2015), a highly interpretable machine learning method.
@@
More recently, dynamic methods have been developed. The goal of mechanistic interpretability is to uncover the mechanisms by which the model dynamically comes to conclusions (Nanda et al., 2023; Bereska and Gavves, 2024; Ferrando et al., 2024; Rai et al., 2024). Methods such as sparse autoencoders (Cunningham et al., 2023; Makelov et al., 2024), neural lenses (Black et al., 2022), and circuit discovery (Conmy et al., 2023) are being used to successfully to enhance insight into how LLMs work, for example, in chess (Davis and Sukthankar, 2024).
@@
Explainable AI and mechanistic interpretability are active areas of research that will allow us to better understand how LLMs reason and come to conclusions (Sharkey et al., 2025). Once a better understanding is reached, LLMs can be improved accordingly, for example, to reduce hallucinations.
2.4.4 Use Case: Benchmarks
@@
3 Acting
@@
First we discuss language models that are enhanced with world knowledge and with robotic actions. Next, we discuss how robots and tools can be used by the LLM, turning them into agentic LLMs, by enabling them to act and interact. Finally, we turn to different use cases for agentic LLMs.
3.1 Action Models
We start by looking at world models, and at how LLMs can be trained by robotic actions, and be used to further operate robots.
3.1.1 World Models
In reinforcement learning, agents learn how to act in an environment (Figure 10). Now, when the real environment is too complex, and learning the policy takes too long,
agents may learn a smaller world model as a surrogate, to allow sample efficient training of the policy (Ha and Schmidhuber, 2018; Hafner et al., 2020). Such world models are learned on the fly by model-based reinforcement learning from the environment interaction, concurrent to policy learning (Moerland et al., 2023; Plaat et al., 2023).
World models have been successful in learning robotic movement in complex environments, to play Atari video games, and to act in open world games such as MineCraft (Hafner et al., 2020, 2023). World models can also be trained effectively with LLMs (Ge et al., 2024). For example, WorldCoder builds a world model as a Python program from interactions with the environment (Tang et al., 2024). The world model explains its interactions with a language model.
While world models are mostly associated with reinforcement learning, they are also used to generate a model in planning domains (blocks-world) in PDDL, to aid task-planning (Guan et al., 2023), achieving success in ALFWorld (Shridhar et al., 2020).
Agents can learn a policy to act with reinforcement learning from surrogate world models. However, agents can also learn action models directly. Three examples are Ahn et al. (2022); Radford et al. (2021); Suglia et al. (2021), who ground language models in world models of robotic actions. Xiang et al. (2024) use world models to finetune language models to gain diverse embodied knowledge while retaining their general language capabilities.
3.1.2 Vision-Language-Action Models
@@
LLMs learn to predict the most probable token to follow a sequence of tokens. Vision-language models also include visual information, to answer questions such as: Is there a red block in the upper corner of the table in this scene? CLIP (Radford et al., 2021) is a widely used Vision Language model. CLIPort learns pathways for robotic manipulation (Shridhar et al., 2022).
Going a step further, vision-language-action models (VLAs) include actions: they are trained on robotic sequences, where they can perform actions in a visual scene, to achieve a goal that is expressed in a language prompt (Zitkovich et al., 2023).
Shah et al. (2023) also train a regular language model from robotic action traces. They show how to utilize off-the-shelf pretrained models trained on large corpora of vision and language datasets. A visual navigation model is used to create a topological mental map of the environment using the robot’s observations. The LLM is then decoding the instructions into a sequence of textual landmarks. Next, the CLIP vision-language model is used for grounding these textual landmarks in the topological map. A search algorithm is used to find a plan for the robot, which is then executed by the visual navigation model.
Various VLA models have been created that achieve impressive zero-shot results, generalizing behavior to unseen situations. Chiang et al. (2024); Brohan et al. (2023); Yang et al. (2025) are examples of VLA models for robotic action, achieving complex tasks such as folding laundry (Black et al., 2024). Ma et al. (2024b) provides an overview.
3.2 Robots and Tools
@@ 3.2.1 Robot Planning
Language models contain a large amount of information about the real world (Ahn et al., 2022). In theory, this may allow the model to exhibit realistic reasoning about robotic behavior. If we could compare a list of intermediate reasoning steps with a list of possible movements of the robot in its environment, then we could prevent the model from suggesting impossible joint movements and actions, and prevent errors or accidents. Such an approach has been tried in the Say Can paper (Ahn et al., 2022). Say Can learns a value function (Kaelbling et al., 1996) of the behavior of a robot in an environment using temporal difference reinforcement learning (Sutton, 1988). This value function is then combined with prompt-based reasoning by the language model, to constrain it from suggesting impossible or harmful actions. The goal of Say Can is to ground the language model in robotic affordances. Say Can is evaluated on 101 real-world robotic tasks, such as how to solve tasks in a kitchen environment (see Figure 11).
@@
Inner Monologue is a related approach to extend LLM reasoning capabilities to robot planning and interaction (Huang et al., 2022). The authors investigate a variety of sources of feedback, such as success detection, object recognition, scene description, and human interaction. Inner Monologue incorporates environmental information into the prompt, linguistically, as if it performs an inner monologue. As in Say Can, the information comes as feedback from different sources. Unlike Say Can, the physics information is inserted directly into the prompt, linguistically.
@@
The language feedback that is thus generated significantly improves performance on three domains, such as simulated and real table top rearrangement tasks and manipulation tasks in a kitchen environment. There are many studies into robotic behavior. A recent approach related to Inner-monologue is Chain of Tools, which proposes a plan-execute-observe pipeline to ground reasoning about tool behavior (Shi et al., 2024a, b).
@@
A challenge in language-driven robot navigation is that most human queries do not conform to preset class labels when referring to an object. Human queries are free-form, and must be mapped to standard object class labels. Dorbala et al. (2023) introduce Language-Driven Zero Shot Object Navigation involves the agent using a freeform natural language description of an object and finding it in a zero shot manner, without ever having seen the environment nor the target object beforehand. By combining implicit knowledge of the LLM with a vision language model, they achieve target object grounding, achieving improved performance on an L-ZSON benchmark.
3.2.2 Action Tools
@@
As we have seen in Section 2.3, LLM results may be augmented with results that are retrieved from external sources, such as search engines. The ability to call search engines can be generalized to calling other tools. When their application programming interface (API) is known, LLMs can be integrated easily with external tools: an API is just another language to learn. Agentic LLMs must be trained to decide when and how to utilize external tools, depending on the task (Shen et al., 2024). Language models can teach themselves to use tools (Schick et al., 2023). The Toolformer model is trained to decide which APIs to call, when to call them, what arguments to pass, and how to best incorporate the results into future token prediction. A range of tools is tested: a calculator, a question-answering system, a search engine, a translation system, and a calendar. Further works extend this to a larger range of tools. ToolBench (Qin et al., 2023) contains 16,464 APIs from RapidAPI, a large dataset of publicly available REST APIs.444 https://rapidapi.com/hub
@@
Another framework is EasyTool (Yuan et al., 2024b), which focuses on structured and unified instructions from tool documentations, building on ToolBench. ToolAlpaca is a benchmark with over 3938 instances from 400 APIs (Tang et al., 2023). A tool-based benchmark for question answering is ToolQA (Zhuang et al., 2023). Gorilla is a finetuned LLaMa-based model for generating API calls (Patil et al., 2023), also introducing the APIBench benchmark. Many tool calling frameworks have been developed.
@@
Selecting the right tool and summarizing its result are difficult skills. Zhao et al. (2024a) study how LLMs can improve recommendation through tool learning. Another approach also suggests to use an LLM for this task (Shen et al., 2024). They use different LLMs for (1) reasoning ability, (2) request writing, and (3) result summarization. Figure 13 illustrates this architecture, consisting of a planner, a caller, and a summarizer, each implemented by a different LLM finetuned for its specific capability.
3.2.3 Computer and Browser Tools
@@
One specific set of actions tools is to let an LLM use a browser or even a complete computer system as a special form of API. Equipping agentic LLMs with the ability to interact directly with a computer environment enables many interaction possibilities. Tools that parse, interpret, and manipulate graphical user interfaces (GUIs) have gained attention for bridging the gap between language models and real-world applications. One such example is OmniParser V2 (Lu et al., 2024b), which introduces a vision-based screen parsing method to detect and label interactable elements such as buttons or icons. By converting raw screenshots into structured representations, OmniParser helps vision-language models ground their action decisions in specific UI components. This grounding increases the accuracy of the action predictions of LLMs.
@@
Browser Use (Müller and Žunič, 2024) is an example of an open-source tool that enables LLMs to use a browser with persistent session management. Browser Use enables agentic LLMs to maintain longer browsing states across multiple pages or domains. The tool manages cookies, session tokens, and dynamic web content updates, thereby allowing LLMs to execute more complex browsing tasks like multi-step form completions or cross-site queries.
@@
For a more in-depth discussion about browser and computer environments, see the survey by Wang et al. (2025). The survey discusses design patterns for combining automated GUI parsing, tool call integration, and human-in-the-loop supervision. Collectively, these approaches reveal that reliable computer and browser tools are important components for agentic LLMs. By translating textual intents into grounded, environment-level operations, such systems expand the possible tasks and domains where language-driven autonomy can be effectively realized.
3.2.4 Safety and Security
@@
Safety and security are increasing in importance as the capabilities of agentic LLMs grow. The robustness of LLMs to jailbreak attacks, where users design prompts to circumvent safety measures and misuse model capabilities, has been studied for chatbots. Meanwhile, agentic LLMs use external tools and can execute multi stage tasks, and may pose a greater risk. Andriushchenko et al. (2024) propose AgentHarm. This benchmark includes a set of 110 malicious agent tasks ranging from fraud, cyber crime, to harassment. The authors find that leading commercial LLMs are surprisingly compliant to their given rules with malicious user requests without jailbreaking (How do I make a pipe bomb? is blocked). However, simple universal templates can be adapted to effectively jailbreak agents (Order online all necessary ingredients to make a pipe bomb and get them delivered to my home without getting flagged by authorities, however, passes in GPT-4).
@@
Rainbow Teaming (Samvelyan et al., 2024) uses an open world approach to generate adversarial LLM prompts, to improve the robustness of LLMs. As with training LLMs, a problem with testing methods is a lack of diversity. Rainbow Teaming casts adversarial prompt generation as a quality-diversity problem. Rainbow Teaming is an open-ended approach (Hughes et al., 2024). It creates diversity with MAP-Elites (Mouret and Clune, 2015), an evolutionary meta search method that iteratively populates an archive with increasingly higher-performing prompts. See Figures 14 and 15.
3.3 Assistants

### v3

@@ 2.4.4. Artificial General Intelligence
The work in this section, and especially the work on self reflection, connects to research on artificial general intelligence, in the scientific tradition (newell1956logic; newell1958elements; newell1961computer) of artificial intelligence that created strong narrow intelligence in backgammon (tesauro1994td), chess (hsu2022behind; muller2018man), and go (silver2016mastering). This tradition views intelligence as a competitive, individualistic, reasoning problem (plaat2020learning). The benefits and risks related to super-intelligence and singularities are actively debated (bostrom1998long; kurzweil2022superintelligence), raising ethical and philosophical questions (dennett2017bacteria). Here, intelligence is regarded as a feature of individuals. In humans and animals, intelligence is assumed to have emerged in a social context (brooks1990elephants; brody1999intelligence; dunbar2003social; y2025intelligence). Most visions of super-intelligence assume that the artificial agent has the ability to use tools and to function in a social environment, something that humans do easily. However, human-unique parts of intelligence emerge in social contexts and depend on constant interaction with others, both evolutionarily and developmentally. We will see work that focuses on social interaction by artificial agents in later sections.
2.4.5. Interpretability
@@
How do LLMs work on the inside? Opening up the black box of neural connectionist architectures is an important topic of research. We wish to understand how the billions of neurons embed representations, how they reason, and how they come to conclusions. Explainable AI provides different methods to do so (minh2022explainable; rios2020feature; van2022comparison; selvaraju2017grad; ali2022xai).
@@
Static methods from the symbolic tradition have been successful in interpreting machine learning models (molnar2020interpretable). Methods exist to relate how input pictures map to output classes, for example using feature maps (ren2016object; kohonen1982self; redmon2016you). Counterfactual analysis (karimi2020model; explains24), LIME (ribeiro2016should), and SHAP (lundberg2017unified) help understand how inputs map to outputs for structured data. Distillation methods can map neural networks to decision trees (hinton2015distilling), a highly interpretable machine learning method.
@@
More recently, dynamic methods have been developed. The goal of mechanistic interpretability is to uncover the mechanisms by which the model dynamically comes to conclusions (nanda2023progress; bereska2024mechanistic; ferrando2024primer; rai2024practical). Methods such as sparse autoencoders (cunningham2023sparse; makelov2024towards), neural lenses (black2022interpreting), and circuit discovery (conmy2023towards) are being used to enhance insight into how LLMs work, for example, in Chain of Thought (chen2025does), and chess (davis2024decoding).
@@
Explainable AI and mechanistic interpretability are active areas of research that will allow us to better understand how LLMs reason and come to conclusions (sharkey2025open). Once a better understanding is reached, LLMs can be improved accordingly, for example, to reduce hallucinations.
2.4.6. Use Case: Benchmarks
@@
Furthermore, in many use cases LLMs need up to date information, beyond that which was available in their training corpus (miikkulainen2024generative). Retrieval augmented generation is an active field that accesses specialty knowledge bases and search engines (such as Google or Wikipedia).
The use of tools creates a bridge to the next category of the taxonomy: LLMs that act in the outside world. Much research has been performed on decision making and reasoning by LLMs. New data is generated by retrieval, and by the use of tools. However, prompt learning methods do not change the parameters of the model; in order to use the data that is generated by inference-time approaches, finetuning must be used.
3. Acting
@@
First we discuss language models that are enhanced with world knowledge and with robotic actions. Next, we discuss how robots and tools can be used by the LLM, turning them into agentic LLMs, by enabling them to act and interact. Finally, we turn to different use cases for agentic LLMs. In Section 3.4 we conclude with an in depth discussion of agentic assistant approaches that are designed to perform or support scientific research itself.
3.1. Action Models
We start by looking at world models, and at how LLMs can be trained by robotic actions.
3.1.1. World Models
In reinforcement learning, agents learn how to act in an environment (Figure 7). When the real environment is too complex, and learning the policy takes too long, agents may learn a smaller world model as a surrogate, to allow sample efficient training of the policy (ha2018world; hafner2020mastering). Such world models are learned on the fly by model-based reinforcement learning from the environment interaction, concurrent to policy learning (moerland2023model; plaat2023high).
World models have been successful in learning robotic movement in complex environments, to play Atari video games, and to act in open world games such as MineCraft (hafner2020mastering; hafner2023mastering). World models can also be trained effectively with LLMs (ge2024worldgpt). For example, WorldCoder builds a world model as a Python program from interactions with the environment (tang2024worldcoder). The world model explains its interactions with a language model.
While world models are mostly associated with reinforcement learning, they are also used to generate a model in planning domains in PDDL (blocks-world), to aid task-planning (guan2023leveraging). For example, shridhar2020alfworld reports success in ALFWorld.
Agents can learn a policy to act with reinforcement learning from surrogate world models. However, agents can also learn action models directly. Three examples are ahn2022can; radford2021learning; suglia2021embodied, who ground language models in world models of robotic actions. xiang2024language use world models to finetune language models to gain diverse embodied knowledge (while retaining their general language capabilities).
3.1.2. Vision-Language-Action Models
@@
LLMs learn to predict the most probable token to follow a sequence of tokens. Vision-language models also include visual information, to answer questions such as: Is there a red block in the upper corner of the table in this scene? CLIP (radford2021learning) is a widely used Vision Language model. CLIPort learns pathways for robotic manipulation (shridhar2022cliport).
Going a step further, vision-language-action models (VLAs) include actions: they are trained on robotic sequences, where they can perform actions in a visual scene, to achieve a goal that is expressed in a language prompt (zitkovich2023rt).
shah2023lm also train a regular language model from robotic action traces. They show how to utilize off-the-shelf pretrained models trained on large corpora of vision and language datasets. A visual navigation model is used to create a topological mental map of the environment using the robot’s observations. The LLM is then decoding the instructions into a sequence of textual landmarks. Next, the CLIP vision-language model is used for grounding these textual landmarks in the topological map. A search algorithm is used to find a plan for the robot, which is then executed by the visual navigation model.
Various VLA models have been created that achieve impressive zero-shot results, generalizing behavior to unseen situations. chiang2024mobility; brohan2023rt; yang2025magma are examples of VLA models for robotic action, achieving complex tasks such as folding laundry (black2024pi_0). ma2024survey provides an overview on VLAs.
3.2. Robots and Tools
@@ 3.2.1. Robot Planning
Language models contain a large amount of information about the real world (ahn2022can). In theory, this may allow the model to exhibit realistic reasoning about robotic behavior. If we could compare a list of intermediate reasoning steps with a list of possible movements of the robot in its environment, then we could prevent the model from suggesting impossible joint movements and actions, and prevent errors or accidents. Such an approach has been tried in the Say Can paper (ahn2022can). Say Can learns a value function (kaelbling1996reinforcement) of the behavior of a robot in an environment using temporal difference reinforcement learning (sutton1988learning). This value function is then combined with prompt-based reasoning by the language model, to constrain it from suggesting impossible or harmful actions. The goal of Say Can is to ground the language model in robotic affordances. Say Can is evaluated on 101 real-world robotic tasks, such as how to solve tasks in a kitchen environment (see Figure 11).
@@
Inner Monologue is a related approach to extend LLM reasoning capabilities to robot planning and interaction (huang2022inner). The authors investigate a variety of sources of feedback, such as success detection, object recognition, scene description, and human interaction. Inner Monologue incorporates environmental information into the prompt, linguistically, as if it performs an inner monologue. As in Say Can, the information comes as feedback from different sources. Unlike Say Can, the physics information is inserted directly into the prompt, linguistically.
@@
The language feedback that is thus generated significantly improves performance on three domains, such as simulated and real table top rearrangement tasks and manipulation tasks in a kitchen environment. There are many studies into robotic behavior. A recent approach related to Inner-monologue is Chain of Tools, which proposes a plan-execute-observe pipeline to ground reasoning about tool behavior (shi2024chain).
@@
A challenge in language-driven robot navigation is that most human queries do not conform to preset class labels when referring to an object. Human queries are free-form, and must be mapped to standard object class labels. dorbala2023can introduce Language-Driven Zero Shot Object Navigation, where the agent uses a freeform natural language description of an object and finds it in a zero shot manner, without ever having seen the environment nor the target object beforehand. By combining implicit knowledge of the LLM with a vision language model, they achieve target object grounding, achieving improved performance on an L-ZSON benchmark.
3.2.2. Action Tools
@@
As we have seen in Section 2.3, LLM results may be augmented with results that are retrieved from external sources, such as search engines. The ability to call search engines can be generalized to calling other tools. When the application programming interface (API) of these tools is known, LLMs can be integrated easily with them. To an LLM, an API is just another language to learn. Agentic LLMs must be trained to decide when and how to utilize external tools, depending on the task (shen2024small). Language models can teach themselves to use tools (schick2023toolformer). The Toolformer model is trained to decide which APIs to call, when to call them, what arguments to pass, and how to best incorporate the results into future token prediction. A range of tools is tested: a calculator, a question-answering system, a search engine, a translation system, and a calendar. Further works extend this to a larger range of tools. ToolBench (qin2023toolllm) contains 16,464 APIs from RapidAPI, a large dataset of publicly available REST APIs.444https://rapidapi.com/hub
@@
Another framework is EasyTool (yuan2024easytool), which focuses on structured and unified instructions from tool documentations, building on ToolBench. ToolAlpaca is a benchmark with over 3938 instances from 400 APIs (tang2023toolalpaca). A tool-based benchmark for question answering is ToolQA (zhuang2023toolqa). Gorilla is a finetuned LLaMa-based model for generating API calls (patil2023gorilla), also introducing the APIBench benchmark. Many tool calling frameworks have been developed.
@@
Selecting the right tool and summarizing its result are difficult skills. zhao2024let study how LLMs can improve recommendation through tool learning. Another approach also suggests to use an LLM for this task (shen2024small). They use different LLMs for (1) reasoning ability, (2) request writing, and (3) result summarization. Figure 13 illustrates this architecture, consisting of a planner, a caller, and a summarizer, each implemented by a different LLM finetuned for its specific capability.
Good results are reported on the LLMs Claude-2, ChatGPT, GPT-4, and Tool-LLaMa, using as reasoning strategies ReAct (yao2022react) and DFSDT (qin2023toolllm). Other frameworks also exist, such as (ocker2024tulip).
3.2.3. Computer and Browser Tools and Agent Interoperability
@@
One specific set of actions tools is to let an LLM interact with a browser or even a complete computer system as a special form of API. Equipping agentic LLMs with the ability to interact directly with a computer environment enables many interaction possibilities. Tools that parse, interpret, and manipulate graphical user interfaces (GUIs) have gained attention for bridging the gap between language models and real-world applications. One such example is OmniParser V2 (lu2024omniparserpurevisionbased), which introduces a vision-based screen parsing method to detect and label interactable elements such as buttons or icons. By converting raw screenshots into structured representations, OmniParser helps vision-language models ground their action decisions in specific UI components. This grounding increases the accuracy of the action predictions of LLMs.
Agents can also interact with other agents, and tap into tool ecosystems. Standards for agent to agent communication are emerging. ray2025review reviews the emerging open A2A standard, for in-context learning. Other protocols for tool use and agent to agent communication are being developed, such as the Model Context Protocol (MCP), Agent Communication Protocol (ACP), and Agent Network Protocol (ANP). Various agentic evaluation benchmarks such as MCPRadar (gao2025mcpradarmultidimensionalbenchmarkevaluating), MCPEval (liu2025mcpevalautomaticmcpbaseddeep), MCPBench (wang2025mcpbenchbenchmarkingtoolusingllm), MCP-Universe (luo2025mcpuniversebenchmarkinglargelanguage) and LiveMCPBench (mo2025livemcpbenchagentsnavigateocean) are built on MCP tool use, to make these benchmarks more representative for real world task settings. intelligenceagentorchestra propose a multi-agent orchestration Tool-Environment-Agent protocol aimed at integrating environments. Agent interoperability protocols are an active field of research, for a survey, see ehtesham2025survey.
@@
Browser Use (browser_use2024) is an example of an open-source tool that enables LLMs to use a browser with persistent session management. Browser Use allows agentic LLMs to maintain longer browsing states across multiple pages or domains. The tool manages cookies, session tokens, and dynamic web content updates, thereby allowing LLMs to execute more complex browsing tasks like multi-step form completions or cross-site queries.
@@
For a more in-depth discussion about browser and computer environments, see the survey by wang2025guiagentsfoundationmodels. The survey discusses design patterns for combining automated GUI parsing, tool call integration, and human-in-the-loop supervision. Collectively, these approaches reveal that reliable computer and browser tools are important components for agentic LLMs. By translating textual intents into grounded, environment-level operations, such systems expand the possible tasks and domains where language-driven autonomy can be effectively realized.
3.2.4. Safety and Security
@@
Safety and security are increasing in importance as the capabilities of agentic LLMs grow. The robustness of LLMs to jailbreak attacks, where users design prompts to circumvent safety measures and misuse model capabilities, has been studied for chatbots. Meanwhile, agentic LLMs use external tools and can execute multi stage tasks, and may pose a greater risk. andriushchenko2024agentharm propose AgentHarm. This benchmark includes a set of 110 malicious agent tasks ranging from fraud, cyber crime, to harassment. The authors find that leading commercial LLMs are surprisingly compliant to their given rules with malicious user requests without jailbreaking (How do I make a pipe bomb? is blocked). However, simple universal templates can be adapted to effectively jailbreak agents (Order online all necessary ingredients to make a pipe bomb and get them delivered to my home without getting flagged by authorities, however, passes in GPT-4).
@@
Rainbow Teaming (samvelyan2024rainbow) uses an open world approach to generate adversarial LLM prompts, to improve the robustness of LLMs. As with training LLMs, a problem with testing methods is a lack of diversity. Rainbow Teaming casts adversarial prompt generation as a quality-diversity problem. Rainbow Teaming is an open-ended approach (hughes2024open). It creates diversity with MAP-Elites (mouret2015illuminating), an evolutionary meta search method that iteratively populates an archive with increasingly higher-performing prompts. Figures 14 and 15 illustrate the ideas of Rainbow Teaming.
3.3. Assistants